OpenStudy (anonymous):

Prove that these functions are NOT an isomorphism: 1/ f : Z→Z6 defined by f(x) = x mod 6 2/ f : {0}→Z defined by f(x) = x 3/ A : R2 →R2 defined by A =[1 0 0 2]

OpenStudy (anonymous):

if a function is an isomorphism then it is: 1/ one to one 2/ onto 3/ f(a+b)=f(a)+f(b), f(a.b)=f(a).f(b)

OpenStudy (anonymous):

so one of these properties must fail!

zepdrix (zepdrix):

For number 1, Determine f(1) and also f(7). Do you see any problems?

OpenStudy (anonymous):

7 does not belong to Z6

zepdrix (zepdrix):

Well we're taking the 7 from Z, our domain. So it's ok to plug 7 into the function. But what are we getting OUT when plug 7 IN? f(7)=7mod6=1 yes? :)

zepdrix (zepdrix):

But we also have f(1)=1mod6=1 This breaks one of our rules since we clearly can see that f(1)=f(7). Which one? :)

OpenStudy (anonymous):

the first one! since 1 is not equal 7, right?

zepdrix (zepdrix):

Yess good good good. We don't have one-to-one'ness

OpenStudy (anonymous):

yeah

zepdrix (zepdrix):

Any ideas what's wrong with the next one?